# SegQuant: A Semantics-Aware and Generalizable Quantization Framework for Diffusion Models

### v3

@@ Appendix B More Experimental Details and Results
SD3.5-DiT 8/8(int) Q-Diffusion 161.22161.22161.22 1.973-1.973- 1.973 0.4600.4600.460 17.9217.9217.92 0.6910.6910.691 7.807.807.80 0.2540.2540.254
PTQD 53.3853.3853.38 0.454-0.454- 0.454 0.4160.4160.416 18.0818.0818.08 0.5820.5820.582 10.0010.0010.00 0.2960.2960.296
PTQ4DiT 15.7115.7115.71 0.4850.4850.485 0.4300.4300.430 18.0818.0818.08 0.4450.4450.445 12.3312.3312.33 0.4920.4920.492
TAC 17.2217.2217.22 0.4300.4300.430 0.4320.4320.432 18.1018.1018.10 0.4610.4610.461 12.1112.1112.11 0.4780.4780.478
SegQuant-A 11.9911.9911.99 0.6390.6390.639 0.4400.4400.440 18.0718.0718.07 0.4070.4070.407 12.7412.7412.74 0.5210.5210.521
SegQuant-G 11.4211.4211.42 0.6390.6390.639 0.4470.4470.447 18.0618.0618.06 0.4120.4120.412 12.4012.4012.40 0.5160.5160.516
4/8(int) PTQ4DiT 67.2367.2367.23 0.314-0.314- 0.314 0.4030.4030.403 18.2118.2118.21 0.5660.5660.566 10.9410.9410.94 0.3860.3860.386
SVDQuant 20.0320.0320.03 0.5180.5180.518 0.4300.4300.430 18.1418.1418.14 0.4460.4460.446 12.3412.3412.34 0.4950.4950.495
SegQuant-G 17.1617.1617.16 0.5760.5760.576 0.4350.4350.435 18.1418.1418.14 0.4380.4380.438 12.2512.2512.25 0.4970.4970.497
SDXL-UNet 8/8(int) Q-Diffusion 64.8764.8764.87 1.466-1.466- 1.466 0.4020.4020.402 17.6517.6517.65 0.5520.5520.552 13.6413.6413.64 0.4480.4480.448
PTQ4DiT 7.157.157.15 0.4720.4720.472 0.4150.4150.415 17.9417.9417.94 0.1970.1970.197 19.6319.6319.63 0.6860.6860.686
SegQuant-A 5.675.675.67 0.4870.4870.487 0.4160.4160.416 17.9417.9417.94 0.1320.1320.132 21.6921.6921.69 0.7520.7520.752
SegQuant-G 5.475.475.47 0.5040.5040.504 0.4160.4160.416 17.9417.9417.94 0.1230.1230.123 22.1122.1122.11 0.7630.7630.763
8/8(fp) Q-Diffusion 5.065.065.06 0.4740.4740.474 0.4060.4060.406 17.8917.8917.89 0.1020.1020.102 23.2823.2823.28 0.7800.7800.780
SegQuant-A 4.964.964.96 0.4750.4750.475 0.4060.4060.406 17.8917.8917.89 0.1010.1010.101 23.3423.3423.34 0.7820.7820.782
SegQuant-G 4.774.774.77 0.4740.4740.474 0.4050.4050.405 17.8917.8917.89 0.0920.0920.092 23.8223.8223.82 0.7930.7930.793
FLUX-DiT 8/8(int) Q-Diffusion 8.968.968.96 0.5130.5130.513 0.4490.4490.449 18.0118.0118.01 0.2980.2980.298 14.4014.4014.40 0.5630.5630.563
PTQ4DiT 13.0413.0413.04 0.4380.4380.438 0.4720.4720.472 17.9717.9717.97 0.5520.5520.552 9.229.229.22 0.3210.3210.321
SegQuant-A 5.725.725.72 0.5920.5920.592 0.4520.4520.452 18.0118.0118.01 0.1490.1490.149 18.2918.2918.29 0.7150.7150.715
SegQuant-G 5.325.325.32 0.5900.5900.590 0.4510.4510.451 18.0218.0218.02 0.1310.1310.131 18.9718.9718.97 0.7370.7370.737
4/8(int) PTQ4DiT 60.2460.2460.24 0.393-0.393- 0.393 0.4180.4180.418 17.8817.8817.88 0.5920.5920.592 10.9810.9810.98 0.4650.4650.465
SVDQuant 8.128.128.12 0.5750.5750.575 0.4420.4420.442 18.0318.0318.03 0.2240.2240.224 16.1916.1916.19 0.6350.6350.635
SegQuant-G 8.568.568.56 0.5660.5660.566 0.4460.4460.446 18.0318.0318.03 0.2190.2190.219 16.3116.3116.31 0.6390.6390.639

### v4

@@ Appendix B More Experimental Details and Results
SD3.5-DiT 8/8(int) Q-Diffusion $161.22$ 1.973-1.973 $0.460$ $17.92$ $0.691$ $7.80$ $0.254$
PTQD $53.38$ 0.454-0.454 $0.416$ $18.08$ $0.582$ $10.00$ $0.296$
PTQ4DiT $15.71$ $0.485$ $0.430$ $18.08$ $0.445$ $12.33$ $0.492$
TAC $17.22$ $0.430$ $0.432$ $18.10$ $0.461$ $12.11$ $0.478$
SegQuant-A $11.99$ $0.639$ $0.440$ $18.07$ $0.407$ $12.74$ $0.521$
SegQuant-G $11.42$ $0.639$ $0.447$ $18.06$ $0.412$ $12.40$ $0.516$
4/8(int) PTQ4DiT $67.23$ 0.314-0.314 $0.403$ $18.21$ $0.566$ $10.94$ $0.386$
SVDQuant $20.03$ $0.518$ $0.430$ $18.14$ $0.446$ $12.34$ $0.495$
SegQuant-G $17.16$ $0.576$ $0.435$ $18.14$ $0.438$ $12.25$ $0.497$
SDXL-UNet 8/8(int) Q-Diffusion $64.87$ 1.466-1.466 $0.402$ $17.65$ $0.552$ $13.64$ $0.448$
PTQ4DiT $7.15$ $0.472$ $0.415$ $17.94$ $0.197$ $19.63$ $0.686$
SegQuant-A $5.67$ $0.487$ $0.416$ $17.94$ $0.132$ $21.69$ $0.752$
SegQuant-G $5.47$ $0.504$ $0.416$ $17.94$ $0.123$ $22.11$ $0.763$
8/8(fp) Q-Diffusion $5.06$ $0.474$ $0.406$ $17.89$ $0.102$ $23.28$ $0.780$
SegQuant-A $4.96$ $0.475$ $0.406$ $17.89$ $0.101$ $23.34$ $0.782$
SegQuant-G $4.77$ $0.474$ $0.405$ $17.89$ $0.092$ $23.82$ $0.793$
FLUX-DiT 8/8(int) Q-Diffusion $8.96$ $0.513$ $0.449$ $18.01$ $0.298$ $14.40$ $0.563$
PTQ4DiT $13.04$ $0.438$ $0.472$ $17.97$ $0.552$ $9.22$ $0.321$
SegQuant-A $5.72$ $0.592$ $0.452$ $18.01$ $0.149$ $18.29$ $0.715$
SegQuant-G $5.32$ $0.590$ $0.451$ $18.02$ $0.131$ $18.97$ $0.737$
4/8(int) PTQ4DiT $60.24$ 0.393-0.393 $0.418$ $17.88$ $0.592$ $10.98$ $0.465$
SVDQuant $8.12$ $0.575$ $0.442$ $18.03$ $0.224$ $16.19$ $0.635$
SegQuant-G $8.56$ $0.566$ $0.446$ $18.03$ $0.219$ $16.31$ $0.639$